If a=i+j + 5k and b =i+j + 2k Compute the cross product a x b. axb= i + j+ k
Answer #1

Given that a= i +J+SK b= i +5+ 2k J axb= k in 2 - (2-5) - 3(2-5) +k(1-1) = 3 i +3J 3. (axb= . Bi +35